OpenStudy (anonymous):

For the graph of y=1/(x-1) what are the horizontal and vertical asymptotes.. I know the vertical asymptote is x=1 and the horizontal one is y= 0 but I do not know how to get these answers

OpenStudy (lgbasallote):

vertical asymptote is the value of x that will make y infinity...i said before that vertical asymptotes occur in rational expressions like this...to find the vertical asymptote just equate your denominator to 0 got it?

OpenStudy (anonymous):

plug in zero for each one so: y-1=(0)-1 and on and y= 0 then do it again with y=0 and x=1

OpenStudy (anonymous):

Yes i understood that for the vertical asymptote.

OpenStudy (anonymous):

horizontal is the same way

OpenStudy (anonymous):

(0)-1=x-1 then 0=x-1 add 1 on both sides and x=1

OpenStudy (anonymous):

The vertical asymptote is when the outcome when an x is plugged in, is nonexistent, the horizontal asymptote is the value that y cannot be. for \[y=1/(1-x)\] when x is 1, 1-1 equals 0 and 1 divided by 0 does not exist, when y equals 0, the expression cannot make 0 because 1/ anything will never give 0, therefore its an asymptote